OpenStudy (anonymous):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points) @itsbribro @sammixboo @thomaster

OpenStudy (itsbribro):

im sorry i dont know

OpenStudy (anonymous):

@dan815 @camerondoherty @Nnesha

OpenStudy (anonymous):

@linn99123

OpenStudy (anonymous):

@Skyz @KlOwNlOvE

OpenStudy (anonymous):

I think it is asking for elimination multipply the first equation by -3 let me know what you get

OpenStudy (anonymous):

-3(-3x + 7y = -16)

OpenStudy (anonymous):

i dont know what you mean. like plug -3 into x and y?

OpenStudy (anonymous):

naw just multiply it by everything

OpenStudy (anonymous):

9x+(-21)y=48

OpenStudy (anonymous):

okay and that negative will be negative 9x-21y=48 -9x+5y=16 now combine like terms by=c

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

combine like terms with the 2 equations

OpenStudy (anonymous):

-16y=64

OpenStudy (anonymous):

then y=-4

OpenStudy (anonymous):

okay now divide 64 by -16

OpenStudy (anonymous):

yes now substitute -4 in for y in one of those equations

OpenStudy (anonymous):

-9x+(-20)=16

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

-9x-20=16 now solve for x

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

@KlOwNlOvE now what

OpenStudy (anonymous):

okay so x=-4 and y=-4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so show how -9x-20=16 is equivalent to -9+5y=16

OpenStudy (anonymous):

you mean -9x correct?

OpenStudy (anonymous):

@camerondoherty @satellite73 can one of yall help him have a feeling i did it the wrong way and gtg

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

idk how they are equivalent

OpenStudy (anonymous):

I'm sure i led you in the wrong direction hopefully one of them will come help not real sure might have to substitute

OpenStudy (anonymous):

they normally dont but okay.

OpenStudy (anonymous):

i dougbt they will join

OpenStudy (anonymous):

how do i do Part B: Ranga helped me on part a

OpenStudy (anonymous):

Part A: -3x + 7y = -16 -9x + 5y = 16 Let us replace (2) by adding (2) to two times (1): 2 times (1): -6x + 14y = -32 Add (2): -9x + 5y = 16 -------------------- -15x + 19y = -16 So the second pair of equations are: -3x + 7y = -16 -15x + 19y = -16

OpenStudy (anonymous):

okay now do what we did to the other system

OpenStudy (anonymous):

what do you mean find x and y?